please help me solve this problem:f(x)= cube root of 32(x+3)^3
Answer by rothauserc(4718)

f(x)= cube root of 32(x+3)^3
take the cube root
f(x) = (32(x+3)^3)^(1/3)
f(x) = (32^(1/3) * (x+3) )
f(x) = 3.174802104 * (x+3)
f(x) = 3.174802104x + 9.524406312